OpenStudy (anonymous):

Given m = 2n + 1, what integer between 0 and m is the inverse of 2 modulo m? Answer in terms of n.

OpenStudy (anonymous):

additive inverse?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

why don't we try it with \(n=5\) so \(m=2\times 5+1=11\)

OpenStudy (anonymous):

what is the additive inverse of \(2\) in this case?

OpenStudy (anonymous):

So the additive inverse would be 6.

OpenStudy (anonymous):

In this case.

OpenStudy (anonymous):

no i don't think so \(2+6=8\)

OpenStudy (anonymous):

wait, can you explain that a bit more?

OpenStudy (anonymous):

yes modulo 11, 11 is the zero what would you add to 2 to get 11?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

yeah 9

OpenStudy (anonymous):

suppose \(n=10\) so \(m=21\) what would you add to \(2\) to get \(21\) ?

OpenStudy (anonymous):

19

OpenStudy (anonymous):

ok now we are ready what would you add to \(2\) to get \(2n+1\) ?

OpenStudy (anonymous):

2n-1

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

It's wrong though... I'm doing the homework problems on aops and it says that the answer is wrong...

OpenStudy (anonymous):

can you help me ganeshie8?

ganeshie8 (ganeshie8):

try multiplicative inverse

ganeshie8 (ganeshie8):

try `n+1`

OpenStudy (anonymous):

It's correct. Can you explain your answer?

ganeshie8 (ganeshie8):

in general you need to solve below to find the multiplicative inverse of \(a\) in \(\mod n \): \(ax \equiv 1 \mod n \)

OpenStudy (anonymous):

Thanks!

ganeshie8 (ganeshie8):

for this particular problem, you have : \(2x = 1 ( \mod 2n+1 )\)

ganeshie8 (ganeshie8):

so the question is : which number you need to multiply to 2, to get 1 ?

OpenStudy (anonymous):

n+1

ganeshie8 (ganeshie8):

yes thats easy to guess here, but if the mod is big then you can use euclid algorithm in reverse to find the inverse